OpenStudy (anonymous):

A falling stone is at a certain instant 160 feet above the ground. 2 seconds later it is only 16 feet above the ground. From what height was it dropped? a) 188 feet b) 186 feet c) 185 feet d) 187 feet e) 183 feet

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

let initial height be x then write the equations for distance traversed by the stone at two time insrances t and t+2 each where by \[x-160= 1/2 g t^2\] and \[x-16 = 1/2 g (t+2)^2\] now solve for x , g is acceleration due to gravity , i dont know it value in ft/sec^2

OpenStudy (anonymous):

thank you!